# Impact of Data Distribution on Fairness Guarantees in Equitable Deep Learning

### v1

@@ 3 Main Results
Let $\mathcal{D}_{a_{i}}\sim\mathcal{N}(\mu_{i},\Sigma_{i})$ be the data distribution for demographic group $a_{i}$, and let $\mathcal{D}\sim\mathcal{N}(\mu,\Sigma)$ be the overall data distribution. Let $f(\cdot)$ be a function that maps input $x$ to a discriminative feature $z$ in a metric space, and let $\ell$ be a loss function bounded by $B$, i.e., $|\ell(z,y)|\leq B$ for all $z,y$. Suppose we have a training set ${(x_{j},y_{j})}{j=1}^{n}$ of size $n$ drawn independently from $\mathcal{D}$. Let $\bar{z}$ be the centroid of the features generated by $f$ on the overall data distribution, and let $\bar{z}_{i}$ be the centroid of the features generated by $f$ on the demographic group $a_{i}$. Then, with probability at least $1-\delta$, the expected loss of $f(\cdot)$ on $\mathcal{D}_{a_{i}}$ is bounded as follows:
$\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]$ $\leq\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]+B\cdot d(\bar{z}_{i},\bar{z})+$
$B\sqrt{\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]-\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]}$
where $d(\cdot,\cdot)$ is a distance function in the metric space.
@@
We start by applying Theorem 3.19 to the discriminative feature space:
$\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]\leq$ $\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]+B|\mu{z_{i}}-\mu_{z}|2+$
$B\sqrt{|\Sigma{z_{i}}-\Sigma_{z}|F}$
where $\mu{z_{i}}$ and $\Sigma_{z_{i}}$ are the mean and covariance matrix of the features generated by $f$ on $\mathcal{D}_{a_{i}}$, and $\mu_{z}$ and $\Sigma_{z}$ are the mean and covariance matrix of the features generated by $f$ on $\mathcal{D}$. By the definition of the centroid, we have $\bar{z}_{i}=\mu{z_{i}}$ and $\bar{z}=\mu_{z}$. Therefore, $|\mu_{z_{i}}-\mu_{z}|_{2}=d(\bar{z}_{i},\bar{z})$. Now, let’s focus on the term $|\Sigma_{z_{i}}-\Sigma_{z}|_{F}$. By the properties of the Frobenius norm and the trace operator [45], we have:
$|\Sigma_{z_{i}}-\Sigma_{z}|F$
$=\sqrt{\text{tr}((\Sigma{z_{i}}-\Sigma_{z})^{T}(\Sigma_{z_{i}}-\Sigma_{z}))}$
$=\sqrt{\text{tr}(\Sigma_{z_{i}}^{2})-2\text{tr}(\Sigma_{z_{i}}\Sigma_{z})+\text{tr}(\Sigma_{z}^{2})}$
$=\sqrt{\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]-2\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}}),z^{\prime}\sim f(\mathcal{D})}[d(z,z^{\prime})]+\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]}$
Using the Cauchy-Schwarz inequality, we can bound the cross-term as follows:
$\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}}),z^{\prime}\sim f(\mathcal{D})}[d(z,z^{\prime})]$ $\leq\sqrt{\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]\cdot\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]}$
$\leq\frac{1}{2}\left(\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]+\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]\right)$
Substituting this bound into the expression for $|\Sigma_{z_{i}}-\Sigma_{z}|_{F}$, we get:
@@
$|\Sigma_{z_{i}}-\Sigma_{z}|F$ $\leq\sqrt{\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]-\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]}$
Combining the results for $|\mu_{z_{i}}-\mu_{z}|2$ and $|\Sigma{z_{i}}-\Sigma_{z}|_{F}$, we obtain the desired bound:
$\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}$ $[\ell(f(x),y)]\leq\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]+B\cdot d(\bar{z}_{i},\bar{z})+$
$B\sqrt{\mathbb{E}_{z\sim f(\mathcal{D}_{a_{i}})}[d^{2}(z,\bar{z}_{i})]-\mathbb{E}_{z\sim f(\mathcal{D})}[d^{2}(z,\bar{z})]}$
This completes the proof. ∎
Remark. Corollary 3.20 demonstrates that the expected loss bound for a demographic group $a_{i}$ depends on three terms: (1) the expected loss on the overall data distribution, (2) the distance between the group-specific feature centroid and the overall feature centroid, and (3) the difference in the average squared distances of the features from their respective centroids. This result highlights the importance of learning discriminative features that are well-clustered around their centroids and have similar distributions across different demographic groups [46]. By minimizing the distance between the group-specific centroids and the overall centroid, as well as reducing the discrepancy in the feature distributions, one can tighten the expected loss bound and achieve more equitable performance.
@@ 4 Experiments
CheXpert is a large dataset of chest X-rays labeled for 14 common chest pathologies from associated radiologist reports. The dataset provides three demographic identity attributes: age, gender, and race [12]. Following the split setting in [47, 48], we select a total of 42,884 patients with 127,118 chest X-ray scans. Among these, 76,205 are used for the training set, 12,673 for the validation set, and 38,240 for the test set. In our experiments, we investigate the detection of pleural effusion across race and gender demographic attributes.
HAM10000, consisting of 10,015 dermatoscopic images, was collected over a span of 20 years from the Department of Dermatology at the Medical University of Vienna, Austria, and a dermatology practice in Queensland, Australia, providing a diverse and representative collection of pigmented skin lesions. Following [49], after filtering out images with missing sensitive attributes, we obtained a refined subset of 9,948 images, and grouped the original seven diagnostic labels into two categories: benign and malignant, to simplify the analysis and facilitate binary classification. For this study, we focus on gender and age as demographic attributes in the context of skin cancer detection.
FairFace is a newly curated dataset consisting of approximately 13,000 images from 3,000 new subjects, combined with a reannotated version of the IJB-C dataset [14], resulting in a total of 152,917 facial images from approximately 6,100 unique identities. The dataset is divided into three subsets: 100,186 images for training, 17,138 images for validation, and 35,593 images for testing. It is comprehensively annotated for protected attributes such as gender and skin color, as well as additional features including age group, eyeglasses, head pose, image source, and face size. In this study, we address glasses detection as a binary classification task, distinguishing images with and without eyeglasses, while focusing on the demographic attributes of age, skin color, and gender.
(a) ViT for AMD Detection on Race
(b) ViT for AMD Detection on Gender
(c) ViT for AMD Detection on Ethnicity
(d) EffNet for AMD Detection on Race
(e) EffNet for AMD Detection on Gender
(f) EffNet for AMD Detection on Ethnicity
@@
Figure 1: Feature distribution and AUC comparison of ViT and EfficientNet for AMD detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision dataset
(a) ViT for DR Detection on Race
(b) ViT for DR Detection on Gender
(c) ViT for DR Detection on Ethnicity
(d) EffNet for DR Detection on Race
(e) EffNet for DR Detection on Gender
(f) EffNet for DR Detection on Ethnicity
Figure 2: Feature distribution and AUC comparison of ViT and EfficientNet for DR detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision dataset
(a) ViT for GL Detection on Race
(b) ViT for GL Detection on Gender
(c) ViT for GL Detection on Ethnicity
(d) EffNet for GL Detection on Race
(e) EffNet for GL Detection on Gender
(f) EffNet for GL Detection on Ethnicity
@@
Figure 3: Feature distribution and AUC comparison of ViT and EfficientNet for Glaucoma (GL) detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision dataset.
@@
Implementation Details. We select two deep learning models with different frameworks as our baseline models: the CNN-based EfficientNet [50] and the Transformer-based ViT [51]. All experiments are conducted on an A100 GPU with 80GB of memory. We initialize both EfficientNet and ViT using pre-trained weights provided by TorchVision. During the fine-tuning phase, EfficientNet is trained for 10 epochs with a learning rate of 1e-4 and a batch size of 10. Similarly, ViT is trained for 10 epochs with a learning rate of 1e-4 and a batch size of 50.
@@
Results. We first trained ViT and EfficientNet models on the training sets of the FairVision dataset for detecting DR, AMD, and Glaucoma. After training, the models were applied to the corresponding test sets, where we obtained the encoded features and predicted labels. Using these outputs, we calculated the Euclidean distance of the features to the mean, representing the empirical feature distribution, and evaluated the Area Under the Curve (AUC) accuracy for both the overall and individual demographic groups. The feature distributions and AUC results for DR, AMD, and Glaucoma detection are illustrated in Fig. 1, Fig. 2, and Fig. 3. From the figures, it is evident that both models achieved reasonable detection performance across all three diseases, with EfficientNet consistently outperforming ViT in terms of accuracy. Additionally, we observed significant disparities in AUC values across different demographic groups, underscoring the importance of studying fairness in AI systems. These disparities are particularly pronounced in the results stratified by race, where the Black group exhibits significantly lower detection accuracy compared to the Asian and White groups across all three diseases.
@@
Furthermore, we extended our experiments to the CheXpert, HAM10000, and FairFace datasets to evaluate model performance on additional medical imaging and facial recognition tasks. Considering the superior performance of EfficientNet on the FairVision dataset compared to ViT, we primarily focused on EfficientNet’s performance across demographic groups in these datasets. The results for CheXpert, HAM10000, and FairFace are presented in Fig. 4, Fig. 5, and Fig. 6, respectively. From the results, it is apparent that detection performance also varies significantly across different demographic groups in these datasets. For instance, in the CheXpert dataset, disparities in AUC are observed when analyzing pleural effusion detection performance across race and gender attributes. Similarly, for HAM10000, detection of skin lesions exhibits performance variations across age and gender groups. In the FairFace dataset, the performance of eyeglass detection also shows notable differences among subgroups defined by age, skin color, and gender.
(a) PE Detection on Race
(b) PE Detection on Gender
@@
Figure 4: Feature distribution and AUC comparison of EfficientNet for Pleural Effusion (PE) detection across two demographic attributes, including Race and Gender, on Chexpert dataset.
(a) SC Detection on Gender
(b) SC Detection on Age
@@
Figure 5: Feature distribution and AUC comparison of EfficientNet for Skin Cancer (SC) detection across two demographic attributes, including Gender and Age, on HAM10000 dataset.
(a) Glasses Detection on Age
(b) Glasses Detection on Skin Color
(c) Glasses Detection on Gender
@@
Figure 6: Feature distribution and AUC comparison of EfficientNet for Glasses detection across three demographic attributes, including Age, Skin Color, and Gender, on FairFace dataset.
@@
Discussion. Our experimental results across multiple datasets demonstrate the relationship between data distributions and fairness guarantees as predicted by Theorem 3.19 and Corollary 3.20. Let’s analyze the results for each dataset:
@@
FairVision Dataset For AMD detection (Fig. 1), both ViT and EfficientNet show significant performance variations across demographic groups. With ViT, the overall feature distribution has a mean of 5.92 and standard deviation of 2.46. The Asian group (mean: 6.26, std: 2.53) and Black group (mean: 6.45, std: 2.65) show larger deviations from the overall distribution compared to the White group (mean: 5.81, std: 2.41). According to Theorem 3.19, this translates to potentially higher expected loss for these groups:
For the Asian group:
$E_{(x,y)\sim D_{Asian}}[\ell(f(x),y)]\leq E_{(x,y)\sim D}[\ell(f(x),y)]+0.34B+0.07B$
For the Black group:
$E_{(x,y)\sim D_{Black}}[\ell(f(x),y)]\leq E_{(x,y)\sim D}[\ell(f(x),y)]+0.53B+0.19B$
@@
Similar patterns emerge for DR detection (Fig. 2), where the Black group shows the largest deviation (mean: 6.66, std: 1.93) from the overall distribution (mean: 6.24, std: 2.30), corresponding to lower AUC values. For Glaucoma detection (Fig. 3), the feature distribution differences are particularly pronounced between racial categories, with the Black group consistently showing the largest deviation and correspondingly lower performance.
@@
HAM10000 Dataset For skin cancer detection (Fig. 5), the age-based disparities are particularly notable. The feature distribution for patients aged ¿60 (mean: 12.43, std: 4.16) differs significantly from the overall distribution (mean: 12.71, std: 4.32), leading to performance disparities that align with our theoretical predictions about the relationship between distribution differences and fairness guarantees.
FairFace Dataset The glasses detection results (Fig. 6) show how multiple demographic attributes interact with fairness guarantees. The feature distributions vary across age groups, skin color, and gender, with the age$>$65 group showing the largest deviation (mean: 10.29, std: 2.74) from the overall distribution (mean: 10.21, std: 2.45). This multi-attribute variation demonstrates the complexity of achieving fairness when dealing with intersectional demographic characteristics.
These empirical results consistently support our theoretical framework, showing that groups with feature distributions deviating significantly from the overall distribution tend to experience lower performance. The relationship between distribution differences and fairness guarantees is particularly evident in:
  1. 1.
    The consistent correlation between feature distribution deviation and AUC performance across all datasets
  2. 2.
    The more pronounced fairness issues in racial categories, where distribution differences tend to be larger
  3. 3.
    The compounding effects when multiple demographic attributes deviate from the overall distribution
The results also validate the practical utility of our theoretical bounds in predicting and understanding fairness disparities.

### v2

@@ 3 Main Results
Let $f(\cdot)$ be a function that maps an input $x$ to a discriminative feature $z=f(x)$. Let $\bar{z}_{i}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[f(x)]$ and $\Sigma_{z_{i}}=\text{Cov}_{(x,y)\sim\mathcal{D}_{a_{i}}}(f(x))$ be the centroid and covariance matrix of the features for demographic group $a_{i}$, respectively. Let $l$ be a loss function that is $L_{l}$-Lipschitz. The expected loss of $f(\cdot)$ on group $a_{i}$ is bounded as follows:
[…]
where $d(\cdot,\cdot)$ is the Euclidean distance between the centroids.
@@
The proof is a direct application of Theorem 3.19 to the feature space representations $z=f(x)$. The theorem provides a bound based on the means ($\mu$) and covariances ($\Sigma$) of the underlying data distributions. We apply this bound to the distributions of the features, $z$.
Substituting the feature space statistics for each group $j$ (i.e., $\mu_{z_{j}}$, $\Sigma_{z_{j}}$) into the bound from Theorem 3.19 gives:
[…]
By definition, the feature centroid $\bar{z}_{i}$ is the mean of the feature distribution for group $i$, so $\bar{z}_{i}=\mu_{z_{i}}$. Therefore, the Euclidean norm of the difference in means, $\mu_{z_{i}}-\mu_{z_{j}}||_{2}$, is equivalent to the distance between the feature centroids, $d(\bar{z}_{i},\bar{z}_{j})$. Substituting this notation gives the final result. ∎
4 Fairness-Aware Regularization
Let $\mathcal{G}$ be the set of $k$ demographic groups. For each group $a_{i}\in\mathcal{G}$, let $f(\cdot)$ be the feature extractor of the model, and $z=f(x)\in\mathbb{R}^{d}$ be the learned representation of input $x$. Let $\bar{z}_{i}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[f(x)]$ and $\Sigma_{z_{i}}=\text{Cov}_{(x,y)\sim\mathcal{D}_{a_{i}}}(f(x))$ be the centroid and covariance matrix of the features for demographic group $a_{i}$, respectively. Let $\pi_{j}>0$ be the proportion of group $a_{j}$ in the overall population, such that $\sum_{j=1}^{k}\pi_{j}=1$.
Motivated by Theorem 3.19 and Corollary 3.20, which bound the expected loss on group $a_{i}$ as:
@@ 4 Fairness-Aware Regularization
$\begin{split}\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]&\leq\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]\\ +L_{l}&\sum_{j=1}^{k}\pi_{j}\left(\|\bar{z}_{i}-\bar{z}_{j}\|_{2}+\sqrt{\|\Sigma_{z_{i}}-\Sigma_{z_{j}}\|_{F}}\right),\end{split}$
we propose a fairness regularization term that directly minimizes the theoretical upper bound on group-wise loss disparity. This term penalizes the pairwise distributional shifts in the learned feature space between all groups:
$\mathcal{R}_{\text{fair}}(f)=\sum_{i\in\mathcal{G}}\sum_{j\in\mathcal{G}}\pi_{j}\left(\underbrace{\|\bar{z}_{i}-\bar{z}_{j}\|_{2}}_{\text{centroid gap}}+\underbrace{\sqrt{\|\Sigma_{z_{i}}-\Sigma_{z_{j}}\|_{F}}}_{\text{covariance gap}}\right)$ (4)
This formulation explicitly encourages the alignment of feature distributions between every pair of demographic groups, thereby minimizing the summation term in the theoretical bound. The overall training objective becomes:
$\min_{f}\left(\frac{1}{n}\sum_{j=1}^{n}\ell(f(x_{j}),y_{j})+\lambda\cdot\mathcal{R}_{\text{fair}}(f)\right)$ (5)
where $\lambda>0$ controls the trade-off between accuracy and fairness.
In practice, the group centroids $\bar{z}_{i}$ and covariances $\Sigma_{z_{i}}$ can be estimated using mini-batch statistics or exponential moving averages during training. The gradients of $\mathcal{R}_{\text{fair}}$ with respect to the parameters of $f$ can be computed via backpropagation, enabling end-to-end optimization.
The regularization is directly derived from Corollary 3.20. By minimizing $\mathcal{R}_{\text{fair}}(f)$, we directly tighten the theoretical upper bound on the expected loss for each demographic group, creating a clear and consistent link between our theoretical analysis and the proposed algorithmic solution.
Refer to caption (a) ViT for AMD Det. on Race Refer to caption (b) ViT for AMD Det. on Gender Refer to caption (c) ViT for AMD Det. on Ethnicity Refer to caption (d) EffNet for AMD Det. on Race Refer to caption (e) EffNet for AMD Det. on Gender Refer to caption (f) EffNet for AMD Det. on Ethnicity Refer to caption (g) ViT-FAR for AMD Det. on Race Refer to caption (h) ViT-FAR for AMD Det. on Gender Refer to caption (i) ViT-FAR for AMD Det. on Ethnicity Refer to caption (j) EffNet-FAR for AMD Det. on Race Refer to caption (k) EffNet-FAR for AMD Det. on Gender Refer to caption (l) EffNet-FAR for AMD Det. on Ethnicity Refer to caption (m) Comparison of ViT and ViT-FAR on AMD Det. Refer to caption (n) Comparison of EffNet and EffNet-FAR on AMD Det.
Figure 1: Feature distribution, AUC performance, and AUC comparison of ViT, EfficientNet, and their fairness-aware regularization (FAR) variants (ViT-FAR and EffNet-FAR) for AMD detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision.
@@ 5 Experiments
CheXpert [16] is a large chest X-ray dataset labeled for 14 pathologies, with demographic attributes (age, gender, race). Following [49, 50], we use 42,884 patients (127,118 scans): 76,205 training, 12,673 validation, and 38,240 test images. We study pleural effusion detection across race and gender.
HAM10000 contains 10,015 dermatoscopic images of pigmented skin lesions collected over 20 years. After filtering missing attributes as in [51], we use 9,948 images, relabeled into benign/malignant categories. We evaluate skin cancer detection with respect to gender and age.
FairFace combines 13,000 new images with a reannotated IJB-C subset [18], yielding 152,917 images from 6,100 identities (100,186 train / 17,138 val / 35,593 test). It includes annotations for gender, skin color, age, and eyeglasses. We treat glasses detection as binary classification, focusing on fairness across age, skin color, and gender.
ACS Income [19] contains 1.66M records from the 2018 U.S. Census ACS, capturing demographic and socioeconomic features. We cast income prediction as binary classification ($\mathdollar 50$k/year) and assess fairness across race and gender.
CivilComments-WILDS includes 450,000 crowdsourced comments labeled for toxicity (majority vote, $\geq$10 workers per comment). With 269,038 training, 45,180 validation, 133,782 testing samples, we frame toxicity detection as binary classification and evaluate fairness across race and gender.
Refer to caption (a) ViT for DR Det. on Race Refer to caption (b) ViT for DR Det. on Gender Refer to caption (c) ViT for DR Det. on Ethnicity Refer to caption (d) EffNet for DR Det. on Race Refer to caption (e) EffNet for DR Det. on Gender Refer to caption (f) EffNet for DR Det. on Ethnicity Refer to caption (g) ViT-FAR for DR Det. on Race Refer to caption (h) ViT-FAR for DR Det. on Gender Refer to caption (i) ViT-FAR for DR Det. on Ethnicity Refer to caption (j) EffNet-FAR for DR Det. on Race Refer to caption (k) EffNet-FAR for DR Det. on Gender Refer to caption (l) EffNet-FAR for DR Det. on Ethnicity Refer to caption (m) Comparison of ViT and ViT-FAR on DR Det. Refer to caption (n) Comparison of EffNet and EffNet-FAR on DR Det.
@@
Figure 2: Feature distribution, AUC performance, and AUC comparison of ViT, EfficientNet, and their fairness-aware regularization (FAR) variants (ViT-FAR and EffNet-FAR) for DR detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision.
Refer to caption (a) ViT for GL Det. on Race Refer to caption (b) ViT for GL Det. on Gender Refer to caption (c) ViT for GL Det. on Ethnicity Refer to caption (d) EffNet for GL Det. on Race Refer to caption (e) EffNet for GL Det. on Gender Refer to caption (f) EffNet for GL Det. on Ethnicity Refer to caption (g) ViT-FAR for GL Det. on Race Refer to caption (h) ViT-FAR for GL Det. on Gender Refer to caption (i) ViT-FAR for GL Det. on Ethnicity Refer to caption (j) EffNet-FAR for GL Det. on Race Refer to caption (k) EffNet-FAR for GL Det. on Gender Refer to caption (l) EffNet-FAR for GL Det. on Ethnicity Refer to caption (m) Comparison of ViT and ViT-FAR on GL Det. Refer to caption (n) Comparison of EffNet and EffNet-FAR on GL Det.
@@
Figure 3: Feature distribution, AUC performance, and AUC comparison of ViT, EfficientNet, and their fairness-aware regularization (FAR) variants (ViT-FAR and EffNet-FAR) for Glaucoma (GL) detection across three demographic attributes, including Race, Gender, and Ethnicity, on FairVision.
@@
Implementation Details. For image datasets, we adopt two baseline models with distinct architectures: the CNN-based EfficientNet [52] and the Transformer-based ViT [53]. For tabular datasets, we employ TabTransformer [54], and for natural language datasets, we use RoBERTa [55]. All experiments are conducted on an NVIDIA A100 GPU with 80 GB of memory. EfficientNet and ViT are initialized with pre-trained weights from TorchVision and fine-tuned for 10 epochs with a learning rate of 1e-4; the batch size is set to 10 for EfficientNet and 50 for ViT. TabTransformer is trained for 50 epochs with a learning rate of 1e-4 and a batch size of 1,000. RoBERTa is fine-tuned for 50 epochs with a learning rate of 1e-5 and a batch size of 64.
@@
Results. We first trained ViT and EfficientNet models on the training sets of the FairVision dataset for detecting DR, AMD, and Glaucoma. After training, the models were applied to the corresponding test sets, where we obtained the encoded features and predicted labels. Using these outputs, we calculated the Euclidean distance of the features to the mean, representing the empirical feature distribution, and evaluated the Area Under the Curve (AUC) for both overall and subgroup performance. The feature distributions and AUC results for DR, AMD, and Glaucoma detection are illustrated in Fig. 1, Fig. 2, and Fig. 3. From the figures, it is evident that both models achieved reasonable detection performance across all three diseases, with EfficientNet consistently outperforming ViT. Additionally, we observed significant disparities in AUC values across demographic groups, underscoring the importance of studying fairness in AI systems. These disparities are particularly pronounced in the results stratified by race, where the Black subgroup exhibits markedly lower detection accuracy compared to the Asian and White subgroups. Notably, incorporating fairness-aware regularization (FAR) improves ES-AUC across demographic groups, demonstrating enhanced balance between subpopulations.
@@
We further extended our experiments to the CheXpert, HAM10000, and FairFace datasets to evaluate model performance on additional medical and natural image tasks. Considering the superior performance of EfficientNet on the FairVision dataset compared to ViT, we primarily focused on EfficientNet’s subgroup performance in these datasets. The results for CheXpert, HAM10000, and FairFace are presented in Fig. 4, Fig. 5, and Fig. 6, respectively. From these results, it is apparent that detection performance also varies significantly across demographic groups. For instance, in the CheXpert dataset, disparities in AUC are observed in pleural effusion detection across race and gender attributes. Similarly, for HAM10000, skin lesion detection shows variations across age and gender groups. In the FairFace dataset, eyeglass detection performance differs among subgroups defined by age, skin tone, and gender. Again, FAR consistently improves ES-AUC, suggesting that fairness-aware regularization mitigates disparities and yields more equitable performance across subgroups.
In addition to image datasets, we further evaluated fairness on tabular and natural language benchmarks. For the ACS Income dataset, we trained TabTransformer to predict whether an individual’s annual income exceeds \$50,000. As shown in Fig. 7, the models achieve high overall accuracy but exhibit pronounced disparities across demographic groups, with the Black subgroup consistently underperforming compared to White and Asian counterparts. For the CivilComments-WILDS dataset, we fine-tuned RoBERTa models for toxic comment detection. The results in Fig. 8 demonstrate strong overall AUC yet reveal substantial subgroup gaps, particularly along race and gender. Importantly, incorporating fairness-aware regularization (FAR) improves ES-AUC across both datasets, indicating better performance balance among demographic groups. These observations confirm that fairness challenges are not confined to medical imaging but also arise in structured tabular data and natural language processing tasks.
Refer to caption (a) EffNet for PE Det. on Race Refer to caption (b) EffNet for PE Det. on Gender Refer to caption (c) EffNet-FAR for PE Det. on Race Refer to caption (d) EffNet-FAR for PE Det. on Gender Refer to caption (e) Comparison of EffNet and EffNet-FAR on PE Det.
@@
Figure 4: Feature distribution, AUC performance, and AUC comparison of EfficientNet and EffNet-FAR for Pleural Effusion (PE) detection across two demographic attributes, including Race and Gender, on Chexpert.
Refer to caption (a) EffNet for SC Det. on Gender Refer to caption (b) EffNet for SC Det. on Age Refer to caption (c) EffNet-FAR for SC Det. on Gender Refer to caption (d) EffNet-FAR for SC Det. on Age Refer to caption (e) Comparison of EffNet and EffNet-FAR on SC Det.
@@
Figure 5: Feature distribution, AUC performance, and AUC comparison of EfficientNet and EffNet-FAR for Skin Cancer (SC) detection across two demographic attributes, including Gender and Age, on HAM10000.
Refer to caption (a) EffNet for Glasses Det. on Age Refer to caption (b) EffNet for Glasses Det. on ST Refer to caption (c) EffNet for Glasses Det. on Gender Refer to caption (d) EffNet-FAR for Glasses Det. on Age Refer to caption (e) EffNet-FAR for Glasses Det. on ST Refer to caption (f) EffNet-FAR for Glasses Det. on Gender Refer to caption (g) Comparison of EffNet and EffNet-FAR on Glasses Det.
@@
Figure 6: Feature distribution, AUC performance, and AUC comparison of EfficientNet and EffNet-FAR for Glasses detection across three demographic attributes, including Age, Skin Tone (ST), and Gender, on FairFace.
Refer to caption (a) TabTrans for Income Prediction on Race Refer to caption (b) TabTrans for Income Prediction on Gender Refer to caption (c) TabTrans-FAR for Income Prediction on Race Refer to caption (d) TabTrans-FAR for Income Prediction on Gender Refer to caption (e) Comparison of TabTrans and TabTrans-FAR on Income Prediction
Figure 7: Feature distribution, AUC performance, and AUC comparison of TabTransformer (TabTrans) and TabTransformer-FAR for Income prediction across two demographic attributes, including Race and Gender, on ACSIncome.
Refer to caption (a) RoBERTa for TC Det. on Race Refer to caption (b) RoBERTa for TC Det. on Gender Refer to caption (c) RoBERTa-FAR for TC Det. on Race Refer to caption (d) RoBERTa-FAR for TC Det. on Gender Refer to caption (e) Comparison of RoBERTa and RoBERTa-FAR on TC Det.
Figure 8: Feature distribution, AUC performance, and AUC comparison of RoBERTa and RoBERTa-FAR for Toxic Comment (TC) detection across two demographic attributes, including Race and Gender, on CivilComments-WILDS.
@@
Discussion. Our experimental results across multiple datasets demonstrate the relationship between data distributions and fairness guarantees, a connection formalized by our theoretical framework, particularly Theorem 3.19 and Corollary 3.20. Let’s analyze the results for each dataset:
@@
FairVision Dataset For AMD detection (Fig. 1), both ViT and EfficientNet show significant performance variations across demographic groups. With ViT, the overall feature distribution has a mean of 5.92 and standard deviation of 2.46. The Asian group (mean: 6.26, std: 2.53) and Black group (mean: 6.45, std: 2.65) show larger deviations from the overall distribution compared to the White group (mean: 5.81, std: 2.41). Our Corollary 3.20 predicts that a group’s excess loss is bounded by a weighted sum of the statistical distances (in both feature centroids and covariances) to all other demographic groups. A significant deviation from the overall distribution’s statistics, as observed empirically for the Asian and especially the Black groups, implies large average distances to the other constituent groups that make up the overall population. The theory, therefore, provides a direct explanation for why these groups with more distant feature distributions are expected to have a higher upper bound on their loss, which is consistent with their observed lower AUC performance.
@@
Similar patterns emerge for DR detection (Fig. 2), where the Black group shows the largest deviation (mean: 6.66, std: 1.93) from the overall distribution (mean: 6.24, std: 2.30), corresponding to lower AUC values. For Glaucoma detection (Fig. 3), the feature distribution differences are particularly pronounced between racial categories, with the Black group consistently showing the largest deviation and correspondingly lower performance. These empirical results strongly support our theoretical framework, which links such performance disparities directly to the distances between group-level feature distributions.
@@
HAM10000 Dataset For skin cancer detection (Fig. 5), the age-based disparities are particularly notable. The feature distribution for patients aged >>60 (mean: 12.43, std: 4.16) differs significantly from the overall distribution (mean: 12.71, std: 4.32), leading to performance disparities that align with our theoretical predictions about the relationship between distribution differences and fairness guarantees.
FairFace Dataset The glasses detection results (Fig. 6) show how multiple demographic attributes interact with fairness guarantees. The feature distributions vary across age groups, skin color, and gender, with the age>>65 group showing the largest deviation (mean: 10.29, std: 2.74) from the overall distribution (mean: 10.21, std: 2.45). This multi-attribute variation demonstrates the complexity of achieving fairness when dealing with intersectional demographic characteristics.
ACS Income Dataset The income prediction results (Fig. 7) reveal pronounced disparities across racial subgroups. The feature distribution for the Black subgroup (mean: 4.62, std: 2.21) deviates considerably from the overall distribution (mean: 4.75, std: 2.18), leading to a lower subgroup AUC. This finding is consistent with our theoretical analysis, which suggests that greater divergence between subgroup and overall feature centroids is linked to weaker fairness guarantees and degraded predictive accuracy.
CivilComments-WILDS Dataset For toxic comment detection (Fig. 8), disparities are evident across both race and gender. The feature distribution for comments mentioning the Black identity (mean: 2.25, std: 10.87) differs from the overall distribution (mean: 2.55, std: 10.20), aligning with observed subgroup performance gaps. These results reinforce our theoretical framework, indicating that larger shifts from the global centroid correspond to amplified fairness risks and lower reliability of subgroup performance.